OpenStudy (candy13106):

In one game, Rafael was up to bat 3 times and made 2 hits. In another game, he was up to bat 5 times with no hits. What percent of the times at bat did Rafael make a hit?

OpenStudy (whpalmer4):

How many at bats? How many hits? Batting percentage = hits / at bats * 100% 1 hit in 4 at bats = 1/4*100% = 25%